At a store 5 bags of chips with 5 containers of dip cost $25 and 2 bags of the same chips with 4 containers of the same dip cost $14. Find the cost of each bag of chips and each container of dip.
Answer by rm29924(97)

let x = cost of chips
let y = cost of dip
first equation
5x + 5y = 25
second equation
2x + 4y = 14
get x by itself in the first equation
5x = 25 - 5y
divide both sides by 5 to get x by itself
x=5-y
now substitue this value for x in the second equation
2(5-y)+4y=14
10-2y+4y=14
10+2y=14
subtract 10 from both sides
2y=4
divide both sides by 2 to get y by itself
y=2
dip costs 2 dollars
plus this value for y back in the first equation
5x+5(2)=25
5x+10=25
5x=15
x=3
answer: dip costs $2.00 a container and chips cost $3.00 a bag
check answer:
5(3)+5(2)=25
15+10=25
25=25